Grade 11General Physics

The ground state energy of hydrogen atom is - 13·6 eV. What are the kinetic and potential energies of the electron in this state ?

tanisha singh
12 Years agoGrade 11

1 Answer

Vimal
9 Years ago
From Bohr’s theory of atomic structureLet K.E. of electrons, K = x,Then P.E. of electron, P = - 2xTotal energy, W = K + P = x - 2x = - xGiven, W = - 13.6 eV.∴∴ x = 13.6 eVwhich gives K = 13.6 eV, P = -27.2 eV